OpenStudy (anonymous):

Which of the following fractions can be used in the conversion of 27 cm3 to the unit m3? 1M/10^-2Cm 1ML/1Cm^3 (10^2M/1M)^3 (1M/10^2Cm)^3

OpenStudy (unklerhaukus):

\[1[\text m]=100[\text{cm}]=10^2[\text{cm}]\] \[1[\text m]^3=(100[\text{cm}])^3=\big(10^2[\text{cm}]\big)^3=10^6[\text{cm}]^3\] \[\frac{1[\text m]^3}{10^6[\text{cm}]^3}=1\] \[27 [\text{cm}]^3\\=27 [\text{cm}]^3\times1\\ =27 [\text{cm}]^3\times\frac{1[\text m]^3}{10^6[\text{cm}]^3}\\ =\frac{27}{10^6}[\text m]^3\]